Use a system of linear equations with two variables and two equations to solve.
A jeep and BMW enter a highway running east-west at the same exit heading in opposite directions. The jeep entered the highway 30 minutes before the BMW did, and traveled 9 mph slower than the BMW. After 3 hours from the time the BMW entered the highway, the cars were 397.5 miles apart. Find the speed of each car, assuming they were driven on cruise control.
*Can someone please explain how to correctly solve the problem. I've tried but keeping the incorrect answers. thanks

Answer by Boreal(15235)

speed of jeep=x, time is t+0.5 in hours or 3.5 hours here
speed of B=x+9, time is 3 hours
distance is speed*time
(x+9)*3+(x)(3.5)=397.5
3x+27+3.5x=397.5
6.5x=370.5
x=57 mph for jeep and travels 199.5 miles in 3.5 hours
x+9=66 mph for B and travels 198 miles in 3 hours